jhonyy9 (jhonyy9):

anonymous samthing new in the world of primes : ,,every thwins generate a new prime" - more understandably - than we assum the members of a twin prime +/- 1 we get always one prime number - for example 5+7=12 +/- 1 = 11-13 so this is twin too - 2. example : 11+13=24 +/-1 = 25 - 23 so 23 is prime - opinion please !!!

jhonyy9 (jhonyy9):

anonymous samthing new in the world of primes : ,,every thwins generate a new prime" - more understandably - than we assum the members of a twin prime +/- 1 we get always one prime number - for example 5+7=12 +/- 1 = 11-13 so this is twin too - 2. example : 11+13=24 +/-1 = 25 - 23 so 23 is prime - opinion please !!! 3 months agoReport Abuse Share Question: This Question is Open 42 anonymous Best Response Medals 0 @ganeshie8 @ParthKohli @Kainui 3 months agoReport Abuse 99 Kainui Best Response Medals 0 all primes are of the form 6n±1 since by exhaustion: 6n+2=2(3n+1) 6n+3=3(2n+1) 6n+4=2(3n+2) 6n+6=6(n+1) Numbers 1 or 5 mod 6 are the only numbers (greater than 2 and 3) which can be prime. 6n+1 6n+5=6(m−1)+5=6m−1 So when you add two twin primes together they're of the form: 6n−1+6n+1=12n=6(2n) which is 1 away from another prime (usually) but not always. So your conjecture is about this but it's not true. If you learn some programming you can test your conjecture! 3 months agoReport Abuse 42 anonymous Best Response Medals 0 thank you @Kainui and what is your opinion about i have posted before - with family tree of primes and of twins using this assum the members and multiplie by 3 +/- 1 and we get new twins - and so continue it - thank you for your answers 3 months agoReport Abuse 40 jhonyy9 Medals 0 why ,,anonymous" when everybody know that this is my posted question - think - about prims ??? 5 minutes agoDelete Type your reply EquationDrawAttach File

jhonyy9 (jhonyy9):

@ganeshie8 do you remember on this ?